OpenStudy (anonymous):

In how many ways can you draw a 6 or a face card from a deck of 52 playing cards?

OpenStudy (asevilla5):

there are four 6 cards and the is 3 face cards from every set . hearts,diamonds,spades and club so there is 12 face cards

OpenStudy (anonymous):

@Asevilla5 the answer is 16

OpenStudy (anonymous):

your answer is gone.. @Loser66

OpenStudy (loser66):

:) I delete it, because it violate the code of the site.

OpenStudy (anonymous):

violate? @Loser66

OpenStudy (anonymous):

if then thats the case, could you at least help me begin to solve the problem.. like how do i start? @Loser66

OpenStudy (loser66):

find P(6) find P(face) add them

OpenStudy (anonymous):

and how do i begin to find p(6) ? @Loser66

OpenStudy (loser66):

P(6) =\(\dfrac{4}{52}\) P(face)=\(\dfrac{12}{52}\) ------------------------ total = \(\dfrac{16}{52}\)